# Game-Theoretic Multiagent Reinforcement Learning

### v4

@@ Abstract
Following the remarkable success of the AlphaGo series, significant advances in multi-agent reinforcement learning (MARL) techniques have been witnessed. MARL corresponds to the learning problem in a multi-agent system in which multiple agents learn simultaneously. It is an interdisciplinary domain with a long history that includes game theory, machine learning, stochastic control, psychology, and optimisation. Although MARL has achieved considerable empirical success in solving real-world games, there is a lack of a self-contained overview in the literature that elaborates the game theoretical foundations of modern MARL methods and summarises the recent advances. In fact, the majority of existing surveys are outdated and do not fully cover the recent developments since 2010. In this work, we provide a monograph on MARL that covers both the fundamentals and the latest developments in the research frontier. The goal of our monograph is to provide a self-contained assessment of the current state-of-the-art MARL techniques from a game theoretical perspective. We expect this work to serve as a stepping stone for both new researchers who are about to enter this fast-growing domain and existing domain experts who want to obtain a panoramic view and identify new directions based on recent advances.
@@ 2.1 Problem Formulation: Markov Decision Process
where $\mathbb{E}^{\pi}$ is the expectation under the probability measure $\mathbb{P}^{\pi}$ over the set of infinitely long state-action trajectories $\tau=(s_{0},a_{0},s_{1},a_{1},...)$ and where $\mathbb{P}^{\pi}$ is induced by state transition probability $P$, the policy $\pi$, the initial state $s$ and initial action $a$ (in the case of the Q-function). The connection between the Q-function and value function is $V^{\pi}(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}[Q^{\pi}(s,a)]$ and $Q^{\pi}=\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[R(s,a,s^{\prime})+V^{\pi}(s^{\prime})]$.
@@ 2.3 Solving Markov Decision Processes
${\displaystyle\underbrace{\hat{Q}(s_{t},a_{t})}_{\text{new value}}\leftarrow\underbrace{\hat{Q}(s_{t},a_{t})}_{\text{old value}}+\underbrace{\alpha}_{\text{learning rate}}\cdot\overbrace{{\bigg{(}}\underbrace{{R_{t}}+{\gamma}\cdot{\max_{a\in\mathbb{A}}\hat{Q}(s_{t+1},a)}}_{\text{temporal difference target}}-\underbrace{\hat{Q}(s_{t},a_{t})}_{\text{old value}}{\bigg{)}}}^{\text{temporal difference error }}}$ (2.7)
@@ 3.1.1 Problem Formulation
  • $N$: the number of agents, $N=1$ degenerates to a single-agent MDP, $N\gg 2$ is referred as many-agent cases in this paper.
@@
  • $R^{i}:\mathbb{S}\times\boldsymbol{\mathbb{A}}\times\mathbb{S}\rightarrow\mathbb{R}$: the reward function that returns a scalar value to the $i-th$ agent for a transition from $(s,\bm{a})$ to $s^{\prime}$. The rewards have absolute values uniformly bounded by $R_{\text{max}}$.
@@
We use the superscript of $(\cdot^{i},\cdot^{-i})$ (for example, $\bm{a}=(a^{i},a^{-i})$), when it is necessary to distinguish between agent $i$ and all other $N-1$ opponents.
@@ 3.1.2 Solving Stochastic Games
  • separable reward state independent transition (SR-SIT) SG: the states and the actions have independent effects on the reward function and the transition function depends on only the joint actions, i.e., $\exists\alpha:\mathbb{S}\rightarrow\mathbb{R},\beta:\boldsymbol{\mathbb{A}}\rightarrow\mathbb{R}$ such that these two conditions hold: $1)\ R^{i}(s,\bm{a})=\alpha(s)+\beta(\bm{a}),\forall i\in\{1,...,N\},\forall s\in\mathbb{S},\forall\bm{a}\in\boldsymbol{\mathbb{A}},$ and $2)\ P(\cdot|s^{\prime},\bm{a})=P(\cdot|s,\bm{a}),\forall\bm{a}\in\boldsymbol{\mathbb{A}},\forall s,s^{\prime}\in\mathbb{S}$.
@@
Figure 3.2: The landscape of different complexity classes. Relevant examples are 1) solving the NE in a two-player zero-sum game, $P$-complete [409], 2) solving the NE in a general-sum game, $PPAD$-hard [128], 3) checking the uniqueness of the NE, $NP$-hard [120], 4) checking whether a pure-strategy NE exists in a stochastic game, $PSPACE$-hard [120], and 5) solving Dec-POMDP, $NEXPTIME$-hard [47].
@@ 3.1.3 Special Types of Stochastic Games
[…]
@@ 3.1.4 Partially Observable Settings
We conclude Section 3 by presenting the relationships between the many different types of POSGs through a Venn diagram in Figure 3.3.
@@ 3.2.1 Problem Formulation
Figure 3.4 shows the game tree of two-player Kuhn poker [299]. In Kuhn poker, the dealer has three cards, a King, Queen, and Jack (King$>$Queen$>$Jack), each player is dealt one card (the orange nodes in Figure 3.4), and the third card is put aside unseen. The game then develops as follows.
@@
  • If player two checks, then the higher card wins $1\$$ from the other player.
@@
  • If player one folds, then player two wins $1\$$ from player one.
  • If player one calls, then the higher card wins $2\$$ from the other player.
@@
  • If player two folds, then player one wins $1\$$ from player two.
  • If player two calls, then the higher card wins $2\$$ from the other player.
@@
  • $N$: the number of players. Some EFGs involve a special player called “chance”, which has a fixed stochastic policy that represents the randomness of the environment. For example, the chance player in Kuhn poker is the dealer, who distributes cards to the players at the beginning.
@@ 3.2.2 Normal-Form Representation
A (simultaneous-move) NFG can be equivalently transformed into an imperfect-information EFG [^13] Note that this transformation is not unique, but they share the same equilibria as the original game. This transformation makes the payoff matrix very large, and the actual scalability of many optimization algorithms for large-scale games is often limited by the large payoff matrix of the game. Some researchers have proposed acceleration methods to improve solving scalability [163]. Moreover, this transformation from NFG to EFG does not hold for perfect-information EFGs. [477] [Chapter 5]. Specifically, since the choices of actions by other agents are unknown to the central agent, this could potentially lead to different histories (triggered by other agents) that can be aggregated into one information state for the central agent.
On the other direction, an imperfect-information EFG can also be transformed into an equivalent NFG in which the pure strategies of each agent $i$ are defined by the Cartesian product $\prod_{S^{i}_{j}\in\mathbb{S}^{i}}\chi(S^{i}_{j})$, which is a complete specification [^14] One subtlety of the pure strategy is that it designates a decision at each choice node, regardless of whether it is possible to reach that node given the other choice nodes. of which action to take at every information state of that agent. In the Kuhn poker example, one pure strategy for player one can be check-bet-check-fold-call-fold; altogether, player one has $642^{6}=$ 64 pure strategies, corresponding to $243\times 2^{3}=$ 24 pure strategies for the chance node and $642^{6}=$ 64 pure strategies for player two. The mixed strategy of each player is then a distribution over all its pure strategies. In this way, the NE in NFG in Eq. (3.6) can still be applied to the EFG, and the NE of an EFG can be solved in two steps: first, convert the EFG into an NFG; second, solve the NE of the induced NFG by means of the Lemke-Howson algorithm [474]. If one further restricts the action space to be state-dependent and adopts the discounted accumulated reward at the terminal node, then the EFG recovers to an SG. While the NE of an EFG can be solved through its equivalent normal form, the computational benefit can be achieved by dealing with the extensive form directly; this motivates the adoption of the sequence-form representation of EFGs.
@@ 3.2.3 Sequence-Form Representation
  • $N$: the number of agents, including the chance node, if any, denoted by $c$.
@@
[…]
@@
[…]
@@ 5.1 Taxonomy of MARL Algorithms
In addition to various levels of local knowledge, MARL algorithms can be classified based on the learning paradigm, as shown in Figure 5.1. For example, the $4$th learning paradigm addresses multiagent problems by building a single-agent controller, which takes the joint information from all agents as inputs and outputs the joint policies for all agents. In this paradigm, agents can exchange any information with any other opponent through the central controller. The information that can be exchanged depends on the assumptions about the level of local knowledge describe in the last paragraph. The $5$th paradigm allows agents to exchange information with other agents only during training; during execution, each agent has to act in a decentralised manner, making decisions based on its own observations only. The $6$th category can be regarded as a particular case of paradigm $5$ in that agents are assumed to be interconnected via a (time-varying) network such that information can still spread across the whole network if agents communicate with their neighbours. The most general case is paradigm $2$, where agents are fully decentralised, with no information exchange of any kind allowed at any time, and each agent executes its own policy. Relaxation of paradigm $2$ yields the $1$st and the $3$rd paradigm, where the agents, although they cannot exchange information, share a single set of policy parameters, or, within a pre-defined group, share a single set of policy parameters.
@@ 6.1 Stochastic Team Games
Figure 6.1: Graphical model of the $level$-$k$ reasoning model [559]. The red part is the equivalent graphical model for the multiagent learning problem. The blue part corresponds to the recursive reasoning steps. Subscript $a_{*}$ stands for the level of thinking, not the time step. The opponent policies are approximated by $\rho^{-i}$. The omitted $level$-00 model considers opponents that are fully randomised. Agent $i$ rolls out the recursive reasoning about opponents in its mind (blue area). In the recursion, agents with higher-level beliefs take the best response to the lower-level agents. The higher-level models conduct all the computations that the lower-level models have done, e.g., the $level$-$2$ model contains the $level$-$1$ model by integrating out $\pi_{0}^{i}(a^{i}|s)$.
@@
One substantial benefit of developing a probabilistic framework for multiagent learning is that it can help model the bounded rationality [487]. Instead of assuming perfect rationality and agents reaching NE, bounded rationality accounts for situations in which rationality is compromised; it can be constrained by either the difficulty of the decision problem or the agents’ own cognitive limitations. One intuitive example is the psychological experiment of the Keynes beauty contest [281], in which all players are asked to guess a number between 00 and $100$ and the winner is the person whose number is closest to the $1/2$ of the average number of all guesses. Readers are recommended to pause here and think about which number you would guess. Although the NE of this game is 00, the majority of people guess a number between $13$ and $25$ [121], which suggests that human beings tend to reason only by $1$-$2$ levels of recursion in strategic games [88], i.e., “I believe how you believe how I believe”.
@@
[…]
@@
[…]
in which a Level-$1$ model is built from the learning agent’s perspective. [210, 476] introduced a Level-00 model where no explicit recursive reasoning is considered.
[…]
@@
$\underbrace{\int_{a_{k-2}^{i}}\Big{[}\rho_{k-1}^{-i}(a_{k-1}^{-i}|s,a_{k-2}^{i})\pi_{k-2}^{i}(a_{k-2}^{i}|s)\Big{]}\mathrm{d}a_{k-2}^{i}}_{\text{opponents of level k-1 best responds to agent $i$ of level k-2}}\bigg{\}}\mathrm{d}a_{k-1}^{-i}.\ \ (\textbf{GR2},\ \text{Level-}K).$ (6.8)
@@ 6.3 Networked Multiagent MDPs
In each iteration, $K$ samples are drawn from $\mathcal{D}$. Since $y^{i}_{k}$ is known only to each agent $i$, Eq. (6.14) becomes a typical consensus optimisation problem (i.e., consensus must be reached for $\theta$) [403]. Multiple effective distributed optimisers can be applied to solve this problem, including the DIGing algorithm [402]. Let $g^{i}(\theta^{i})=\frac{1}{2K}\sum_{j=1}^{K}\big{[}y^{i}_{k}-f(s_{k},\bm{a}_{k};\theta)\big{]}^{2}$, $\alpha$ be the learning rate, and $G([N],E_{l})$ be the topology of the network in the $l$st iteration; the DIGing algorithm designs the gradient updates for each agent $i$ as
@@ 6.4 Stochastic Potential Games
[…]
@@ 7.1 Minimax Optimization
$z_{t}=\arg\min_{z\in\mathcal{Z}}\{\eta\langle z,F(z_{t-1})\rangle+D_{\psi}(z,\widehat{z_{t}})\}$ (7.22)
$\widehat{z_{t+1}}=\arg\min_{z\in\mathcal{Z}}\{\eta\langle z,F(z_{t})\rangle+D_{\psi}(z,\widehat{z_{t}})\}$ (7.23)
where $z_{t}=(x_{t},y_{t}),F(z_{t})=(\frac{df(x_{t},y_{t})}{dx},-\frac{df(x_{t},y_{t})}{dy}),\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, $D_{\psi}(u,v)$ is the Bregman divergence in mirror descent. To see that OMDA is a unification of OGDA and OMWU, it can be shown that $D_{\psi}(u,v)=\frac{1}{2}||u-v||^{2}$ for OGDA and $D_{\psi}(u,v)=\text{KL}(u,v)$ for OMWU. The readers can try to prove that by specifying different Bregman divergence, the update rules in Eq. (7.22) and (7.23) will recover the rules of OMDA and OMWU.

### v5

@@ Abstract
Tremendous advances have been made in multiagent reinforcement learning (MARL). MARL corresponds to the learning problem in a multiagent system in which multiple agents learn simultaneously. It is an interdisciplinary field of study with a long history that includes game theory, machine learning, stochastic control, psychology, and optimization. Despite great successes in MARL, there is a lack of a self-contained overview of the literature that covers game-theoretic foundations of modern MARL methods and summarizes the recent advances. The majority of existing surveys are outdated and do not fully cover the recent developments since 2010. In this work, we provide a monograph on MARL that covers both the fundamentals and the latest developments on the research frontier. The goal of this monograph is to provide a self-contained assessment of the current state-of-the-art MARL techniques from a game-theoretic perspective. We expect this work to serve as a stepping stone for both new researchers who are about to enter this fast-growing field and experts in the field who want to obtain a panoramic view and identify new directions based on recent advances.
@@
jun.wang@cs.ucl.ac.uk and Tuomas Sandholm
Carnegie Mellon University, USA
@@ 2.1 Problem Formulation: Markov Decision Process
[…]
@@ 2.3 Solving Markov Decision Processes
[…]
@@ 3.1.1 Problem Formulation
  • NN: the number of agents, $N=1$ degenerates to a single-agent MDP, $N\gg 2$ is referred as many-agent cases in this paper.
@@
  • $R^{i}:\mathbb{S}\times\boldsymbol{\mathbb{A}}\times\mathbb{S}\rightarrow\mathbb{R}$: the reward function that returns a scalar value to the ithi-th agent for a transition from $(s,\bm{a})$ to $s^{\prime}$. The rewards have absolute values uniformly bounded by $R_{\text{max}}$.
@@
We use the superscript of $(\cdot^{i},\cdot^{-i})$ (for example, $\bm{a}=(a^{i},a^{-i})$), when it is necessary to distinguish between agent $i$ and all other N1N-1 opponents.
@@ 3.1.2 Solving Stochastic Games
[…]
@@
Figure 3.2: The landscape of different complexity classes. Relevant examples are 1) solving the NE in a two-player zero-sum game, $P$-complete [409], 2) solving the NE in a general-sum game, PPADPPAD-hard [128], 3) checking the uniqueness of the NE, NPNP-hard [120], 4) checking whether a pure-strategy NE exists in a stochastic game, PSPACEPSPACE-hard [120], and 5) solving Dec-POMDP, NEXPTIMENEXPTIME-hard [47].
@@ 3.1.3 Special Types of Stochastic Games
$\textbf{Normal-form game solver}+\textbf{MDP solver}=\textbf{SG solver},$
@@ 3.1.4 Partially Observable Settings
We conclude Section 3 by presenting the relationships between the many different types of POSGs through a Venn diagram in Figure $3.$
@@ 3.2.1 Problem Formulation
Figure 3.4 shows the game tree of two-player Kuhn poker [299]. In Kuhn poker, the dealer has three cards, a King, Queen, and Jack (King>>Queen>>Jack), each player is dealt one card (the orange nodes in Figure 3.4), and the third card is put aside unseen. The game then develops as follows.
@@
  • If player two checks, then the higher card wins 1\$1\\$ from the other player.
@@
  • If player one folds, then player two wins 1\$1\\$ from player one.
  • If player one calls, then the higher card wins 2\$2\\$ from the other player.
@@
  • If player two folds, then player one wins 1\$1\\$ from player two.
  • If player two calls, then the higher card wins 2\$2\\$ from the other player.
@@
  • NN: the number of players. Some EFGs involve a special player called “chance”, which has a fixed stochastic policy that represents the randomness of the environment. For example, the chance player in Kuhn poker is the dealer, who distributes cards to the players at the beginning.
@@ 3.2.2 Normal-Form Representation
A (simultaneous-move) NFG can be equivalently transformed into an imperfect-information EFG [^13] This transformation is not unique, but they share the same equilibria as the original game. This transformation makes the payoff matrix very large, and the actual scalability of many optimization algorithms for large-scale games is often limited by the large payoff matrix of the game. Some researchers have proposed acceleration methods to improve solving scalability [163]. Moreover, this transformation from NFG to EFG does not hold for perfect-information EFGs. [477] [Chapter 5]. Specifically, since the choices of actions by other agents are unknown to the central agent, this could potentially lead to different histories (triggered by other agents) that can be aggregated into one information state for the central agent.
In the other direction, an imperfect-information EFG can also be transformed into an equivalent NFG in which the pure strategies of each agent $i$ are defined by the Cartesian product $\prod_{S^{i}_{j}\in\mathbb{S}^{i}}\chi(S^{i}_{j})$, which is a complete specification [^14] One subtlety of the pure strategy is that it designates a decision at each choice node, regardless of whether it is possible to reach that node given the other choice nodes. of which action to take at every information state of that agent. In the Kuhn poker example, one pure strategy for player one can be check-bet-check-fold-call-fold; altogether, player one has $2^{6}=64$ pure strategies, corresponding to $3\times 2^{3}=24$ pure strategies for the chance node and $2^{6}=64$ pure strategies for player two. The mixed strategy of each player is then a distribution over all its pure strategies. In this way, the NE in NFG in Eq. (3.6) can still be applied to the EFG, and the NE of an EFG can be solved in two steps: first, convert the EFG into an NFG; second, solve the NE of the induced NFG by means of the Lemke-Howson algorithm [474]. If one further restricts the action space to be state-dependent and adopts the discounted accumulated reward at the terminal node, then the EFG recovers to an SG. While the NE of an EFG can be solved through its equivalent normal form, the computational benefit can be achieved by dealing with the extensive form directly; this motivates the adoption of the sequence-form representation of EFGs.
@@ 3.2.3 Sequence-Form Representation
  • NN: the number of agents, including the chance node, if any, denoted by $c$.
@@
    Note that other players’ actions are not part of agent $i$’s sequence. In the example of Kuhn poker, $\Sigma^{c}=\{\varnothing,\$Jack, Queen, King, Jack-Queen, Jack-King, Queen-Jack, Queen-King, King-Jack, King-Queen}\}, $\Sigma^{1}=\{\varnothing,$ check, bet, check-fold, check-bet}\}, and $\Sigma^{2}=\{\varnothing,\ \text{check, bet, fold, call}\}$.
@@
    The first constraint requires that $\mu^{\pi^{i}}$ is a proper probability distribution. In addition, the second constraint in Eq. (3.15) indicates that in order for a realization plan to be valid to recover a behavioral strategy, at each information state of agent $i$, the probability of reaching that information state must equal the summation of the realization probabilities of all the extended sequences. In the example of Kuhn poker, we have $C^{1}$ for player one by $\mu^{\pi^{1}}(\text{check})=\mu^{\pi^{1}}(\text{check-fold})+\mu^{\pi^{1}}(\text{check-call})$.
@@ 5.1 Taxonomy of MARL Algorithms
In addition to various levels of local knowledge, MARL algorithms can be classified based on the learning paradigm, as shown in Figure 5.1. For example, the $4$th learning paradigm addresses multiagent problems by building a single-agent controller, which takes the joint information from all agents as inputs and outputs the joint policies for all agents. In this paradigm, agents can exchange any information with any other opponent through the central controller. The information that can be exchanged depends on the assumptions about the level of local knowledge describe in the last paragraph. The $5$th paradigm allows agents to exchange information with other agents only during training; during execution, each agent has to act in a decentralised manner, making decisions based on its own observations only. The $6$th category can be regarded as a particular case of paradigm $5$ in that agents are assumed to be interconnected via a (time-varying) network such that information can still spread across the whole network if agents communicate with their neighbours. The most general case is paradigm $2$, where agents are fully decentralised, with no information exchange of any kind allowed at any time, and each agent executes its own policy. Relaxation of paradigm $2$ yields the 11st and the 33rd paradigm, where the agents, although they cannot exchange information, share a single set of policy parameters, or, within a pre-defined group, share a single set of policy parameters.
@@ 6.1 Stochastic Team Games
Figure 6.1: Graphical model of the levellevel-$k$ reasoning model [559]. The red part is the equivalent graphical model for the multiagent learning problem. The blue part corresponds to the recursive reasoning steps. Subscript $a_{*}$ stands for the level of thinking, not the time step. The opponent policies are approximated by $\rho^{-i}$. The omitted levellevel-0 model considers opponents that are fully randomised. Agent $i$ rolls out the recursive reasoning about opponents in its mind (blue area). In the recursion, agents with higher-level beliefs take the best response to the lower-level agents. The higher-level models conduct all the computations that the lower-level models have done, e.g., the levellevel-22 model contains the levellevel-11 model by integrating out $\pi_{0}^{i}(a^{i}|s)$.
@@
One substantial benefit of developing a probabilistic framework for multiagent learning is that it can help model the bounded rationality [487]. Instead of assuming perfect rationality and agents reaching NE, bounded rationality accounts for situations in which rationality is compromised; it can be constrained by either the difficulty of the decision problem or the agents’ own cognitive limitations. One intuitive example is the psychological experiment of the Keynes beauty contest [281], in which all players are asked to guess a number between 0 and $100$ and the winner is the person whose number is closest to the $1/2$ of the average number of all guesses. Readers are recommended to pause here and think about which number you would guess. Although the NE of this game is 0, the majority of people guess a number between $13$ and $25$ [121], which suggests that human beings tend to reason only by 11-22 levels of recursion in strategic games [88], i.e., “I believe how you believe how I believe”.
@@
$\bm{\pi}_{\theta}\left(a^{i},a^{-i}|s\right)=\pi_{\theta^{i}}^{i}\left(a^{i}|s\right)\rho_{\theta^{-i}}^{-i}\left(a^{-i}|s,a^{i}\right)\ \ \ \ (\textbf{PR2},\ \ \text{Level-}1),$ (6.5)
@@
$\bm{\pi}_{\theta}\left(a^{i},a^{-i}|s\right)=\pi_{\theta^{i}}^{i}\left(a^{i}|s,a^{-i}\right)\rho_{\theta^{-i}}^{-i}\left(a^{-i}|s\right)\ \ \ \ (\textbf{ROMMEO},\ \ \text{Level-}1),$ (6.6)
in which a Level-$1$ model is built from the learning agent’s perspective. [210, 476] introduced a Level-0 model where no explicit recursive reasoning is considered.
$\bm{\pi}_{\theta}\left(a^{i},a^{-i}|s\right)=\pi_{\theta^{i}}^{i}\left(a^{i}|s\right)\rho_{\theta^{-i}}^{-i}\left(a^{-i}|s\right)\ \ \ \ (\text{Level-}0).$ (6.7)
@@
[…]
@@ 6.3 Networked Multiagent MDPs
In each iteration, $K$ samples are drawn from $\mathcal{D}$. Since $y^{i}_{k}$ is known only to each agent $i$, Eq. (6.14) becomes a typical consensus optimisation problem (i.e., consensus must be reached for $\theta$) [403]. Multiple effective distributed optimisers can be applied to solve this problem, including the DIGing algorithm [402]. Let $g^{i}(\theta^{i})=\frac{1}{2K}\sum_{j=1}^{K}\big{[}y^{i}_{k}-f(s_{k},\bm{a}_{k};\theta)\big{]}^{2}$, $\alpha$ be the learning rate, and $G([N],E_{l})$ be the topology of the network in the llst iteration; the DIGing algorithm designs the gradient updates for each agent $i$ as
@@ 6.4 Stochastic Potential Games
$\text{s.t.}\quad\sum_{a^{2}\in\mathbb{A}^{2}}R^{1}(a^{1},a^{2})\cdot\pi^{2}(a^{2})$ $\leq U_{1}^{*},\forall a^{1}\in\mathbb{A}^{1}$ (7.2)
@@ 7.1 Minimax Optimization
[…]